### v1

# AAVENUE: Detecting LLM Biases on NLU Tasks in AAVE via a Novel Benchmark

Abstract
Detecting biases in natural language understanding (NLU) for African American Vernacular English (AAVE) is crucial to developing inclusive natural language processing (NLP) systems. To address dialect-induced performance discrepancies, we introduce AAVENUE (AAVE Natural Language Understanding Evaluation), a benchmark for evaluating large language model (LLM) performance on NLU tasks in AAVE and Standard American English (SAE). AAVENUE builds upon and extends existing benchmarks like VALUE, replacing deterministic syntactic and morphological transformations with a more flexible methodology leveraging LLM-based translation with few-shot prompting, improving performance across our evaluation metrics when translating key tasks from the GLUE and SuperGLUE benchmarks. We compare AAVENUE and VALUE translations using five popular LLMs and a comprehensive set of metrics including fluency, BARTScore, quality, coherence, and understandability. Additionally, we recruit fluent AAVE speakers to validate our translations for authenticity. Our evaluations reveal that LLMs consistently perform better on SAE tasks than AAVE-translated versions, underscoring inherent biases and highlighting the need for more inclusive NLP models. We have open-sourced our source code on GitHub and created a website to showcase our work at https://aavenue.live.
AAVENUE: Detecting LLM Biases on NLU Tasks in AAVE via a Novel Benchmark
Abhay Guptathanks: Lead Author           Philip Meng           Ece Yurtseven           Sean O’Brien thanks: Senior Author          Kevin Zhu22footnotemark: 2 Algoverse AI Research abhaygupta1266@gmail.com, kevin@algoverse.us
1 Introduction
NLP systems have shown exceptional performance on various benchmarks, excelling in tasks such as sentiment analysis, machine translation, and question answering (Radford et al., 2019; Brown et al., 2020; OpenAI, 2024; Bubeck et al., 2023). However, these benchmarks feature mainly SAE, often neglecting nonstandard dialects such as AAVE (Blodgett et al., 2020; Weidinger et al., 2021; Sap et al., 2019; Deas et al., 2023). This oversight results in biased and inequitable NLP systems that do not adequately serve a significant portion of speakers.
The advent of LLMs such as GPT-4 has driven advances in NLU tasks, achieving state-of-the-art results across various applications (Radford et al., 2019; Brown et al., 2020; OpenAI, 2024; Bubeck et al., 2023). Despite these advancements, LLMs exhibit persistent biases against nonstandard dialects, including AAVE, particularly in tasks involving natural language generation and toxicity detection (Zhou et al., 2021). These biases highlight the need for comprehensive benchmarks that evaluate model performance in different dialects, ensuring fair and reliable language technologies for all users (Gehrmann et al., 2021).
Benchmarks such as GLUE and SuperGLUE have contributed significantly to the evaluation of NLP models, yet they focus mainly on SAE, neglecting crucial dialectal variations (Wang et al., 2019, 2020). Although VALUE has sought to bridge this gap by using deterministic linguistic transformations to assess model performance in AAVE, these transformations are often context-specific and difficult to generalize, limiting their broader applicability (Ziems et al., 2022, 2023).
To address these limitations, we introduce AAVENUE, a benchmark specifically designed to evaluate LLM performance across dialects, with a particular focus on AAVE. Our goal is to promote inclusivity and equity in NLP tools by providing a more comprehensive assessment of dialectal fairness.
Our contributions are as follows:
  1. 1.
    Creation of Benchmark: We developed AAVENUE, a benchmark to evaluate LLMs on NLU tasks in both AAVE and SAE, using GPT-4o-mini for few-shot prompting. Our translations were validated for cultural and linguistic authenticity and outperformed or were comparable to those from the VALUE benchmark across various metrics and five popular LLMs.
  2. 2.
    Bias Demonstration: Our evaluations of popular LLMs on our benchmark revealed biases, with SAE versions consistently achieving higher accuracy than AAVE translations across key tasks, highlighting the need for more inclusive NLP models.
2 Dataset
To evaluate the performance of models on SAE and AAVE, we selected five key tasks from the GLUE and SuperGLUE benchmarks, knwon for their diversity and relevance to natural language understanding tasks (Wang et al., 2019, 2020). These tasks include BoolQ, MultiRC, SST-2, COPA, and WSC.
2.1 Task Selection
Each task was chosen for its ability to challenge models in different aspects of natural language understanding:
  • BoolQ: BoolQ (Boolean Questions) tests models’ ability to understand and respond to yes/no questions based on a passage, which helps evaluate the model’s ability to process and interpret information across dialects.
  • MultiRC: MultiRC (Multiple Sentence Reading Comprehension) involves answering questions that require understanding and connecting information from different parts of a passage, which tests how well models can handle more complex and interconnected texts in AAVE.
  • SST-2: SST-2 (Stanford Sentiment Treebank) is used to evaluate sentiment analysis, providing insights into whether models understand sentiment differently in SAE compared to AAVE.
  • COPA: COPA (Choice of Plausible Alternatives) challenges models to choose the most likely outcome or cause from two options, focusing on their ability to reason through cause-and-effect scenarios in AAVE.
  • WSC: WSC (Winograd Schema Challenge) tests how well models can determine which noun a pronoun refers to in tricky situations, which is important for understanding different dialects.
2.2 Translating SAE to AAVE
For each task, we sampled 1000 data points and few-shot prompted GPT-4o-mini to translate each data point from SAE to AAVE. We used few-shot examplars sourced from the VALUE benchmark, consisting of AAVE translations that were hand-validated by fluent AAVE speakers. To see the few-shot prompt translations we used, please refer to Table 3 in the Appendix.
2.3 Validation Steps
We assessed the quality of our AAVE translations using a set of carefully chosen metrics: fluency, coherence, understandability, quality, and BARTScores. Fluency measured whether the generated text was well-written and grammatical, with scores out of 100. Coherence evaluated how much the generated text made sense, checking the logical flow and consistency of ideas within the translations, also scored out of 100. Understandability assessed how easily the translation could be comprehended by readers, ensuring that the text is understandable, with scores out of 100. Quality provided an overall assessment of quality of the text. This is also scored out of 100 as well. BARTScores were used to evaluate how closely the translations aligned with the original SAE sentences, with scores closer to 0 indicating better alignment and accuracy. All these metrics were specifically calculated to compare our scores against those from the VALUE benchmark, allowing us to directly evaluate the performance of our translations relative to an established standard.
We also performed comparison scores by zero-shot prompting five large language models to choose between our GPT-4o-mini translations and those from the VALUE benchmark in a binary task. This provided a direct comparison of translation effectiveness.
Finally, we recruited ten fluent AAVE speakers from the Bronx and Queens, NY, to rate the translations on a scale of 1 to 10, focusing on how well they reflect AAVE. The human evaluations ensured our translations accurately matched the AAVE dialect.
2.3.1 Translation Metrics Analysis
The evaluation of our AAVE translations against those from the VALUE benchmark demonstrates clear advantages across several key metrics. (1) Quality: Our translations scored higher in quality, with our BoolQ passages receiving a score of 76.57 compared to VALUE’s 58.21. This shows that our translations are better in terms of overall quality, including accuracy, style, and appropriateness. (see Table 8). (2) Fluency: Our translations achieved a fluency score of 70.72 in BoolQ passages, compared to VALUE’s 57.32. This suggests that our translations are better written and more grammatically sound, resulting in improved readability. (see Table 6). (3) Coherence: Our translations exhibited better coherence, with scores like 74.39 for BoolQ passages versus VALUE’s 62.10, showing that our translations make more logical sense and maintain consistency throughout the text (see Table 7). (4) Understandability: In terms of understandability, our translations were clearer and more accessible, scoring 76.53 in BoolQ passages compared to 52.26 for VALUE. This indicates that our translations are easier for readers to understand (see Table 9). (5) BARTScores: Finally, our translations achieved better BARTScores, with a score of -1.44 in BoolQ passages compared to VALUE’s -1.54, indicating that our translations are closer to human-produced texts and better maintain fidelity to the original content (see Table 10). Collectively, these metrics confirm that our translations consistently outperform those from the VALUE benchmark, providing superior quality, fluency, coherence, understandability, and fidelity.
2.3.2 Comparison Scores Analysis
The comparison scores provide a direct evaluation of our AAVE translations against those from the VALUE benchmark across various tasks and models. As shown in Tables 4 and 5, our AAVE translations were consistently preferred over the VALUE translations by the LLMs we evaluated. For instance, in the BoolQ passage task using GPT-4-turbo, our translations were preferred 94.51% of the time compared to VALUE’s 4.62% (see Table 4). Similarly, in the COPA premise task, GPT-4o-mini showed a preference for our translations 90.42% of the time, with VALUE translations receiving only 9.38% (see Table 4). When evaluated with the Gemini-1.5-flash model on the BoolQ passage task, our translations were preferred 93.85% of the time compared to VALUE’s 6.15% (see Table 5). These consistent preferences across multiple models and tasks show the effectiveness of our translations in capturing the nuances of AAVE while maintaining the meaning of the original text. The results indicate that our approach to translating SAE to AAVE not only meets but often exceeds the standards set by the VALUE benchmark.
2.4 Validation by AAVE Speakers
We recruited 10 fluent AAVE speakers from the Bronx and Queens area to assess the quality of our AAVE translations across five tasks on a scale of 1 to 10, specifically rating how accurately the translations reflect AAVE. The results, detailed in Tables 14 and 15, show that the average scores across the tasks ranged from 7.02 to 7.27. The highest average score was observed for MultiRC (7.27), while BoolQ received the lowest average score (7.02). These consistent ratings across multiple fluent speakers, reinforce the reliability of our AAVE translations. The slightly higher scores for subsets like WSC and MultiRC suggest that our translations were particularly effective in maintaining clarity and coherence in more complex linguistic structures. Overall, the validators’ feedback confirms the quality and authenticity of our translations, aligning well with our evaluation metrics and further validating our approach.
3 Results
3.1 Accuracy Scores Across Tasks
We evaluated the accuracy of the translations in five tasks using five LLMs. The accuracy scores show the performance of each model in SAE and AAVE translations, as shown in Tables 1 and 2.
Table 1: Accuracy Scores for GPT Models (%)
All scores are presented in the format SAE/AAVE.
Task GPT-4o-mini GPT-4-turbo GPT-4o
SST-2 90.40/88.40 (-2.0) 94.00/92.80 (-1.2) 88.80/87.30 (-1.5)
BoolQ 88.29/85.29 (-3.0) 88.09/86.49 (-1.6) 89.19/86.89 (-2.3)
COPA 95.40/93.20 (-2.2) 97.60/96.80 (-0.8) 97.20/96.40 (-0.8)
WSC 60.03/57.90 (-2.1) 69.60/68.69 (-0.9) 70.36/67.02 (-3.3)
MultiRC 84.50/72.00 (-12.5) 86.20/73.70 (-12.5) 87.50/71.30 (-16.2)
Table 2: Accuracy Scores for Gemini Models (%)
All scores are presented in the format SAE/AAVE.
Task Gemini-1.5-Flash Gemini-1.5-Pro
SST-2 87.70/87.10 (-0.6) 92.00/91.40 (-0.6)
BoolQ 89.69/87.29 (-2.4) 89.49/85.89 (-3.6)
COPA 91.40/92.00 (+0.6) 97.40/95.80 (-1.6)
WSC 48.78/48.48 (-0.3) 51.37/51.22 (-0.2)
MultiRC 84.10/70.70 (-13.4) 85.90/71.90 (-14.0)
3.2 Accuracy Score Analysis
The accuracy scores from Tables 1 and 2 highlight consistent performance drops when handling AAVE translations across all models. MultiRC and WSC exhibited the largest declines, indicating challenges in complex reading comprehension and pronoun resolution tasks. GPT-4-turbo generally showed smaller accuracy drops, suggesting better adaptation to AAVE, while other models like GPT-4o-mini struggled more, particularly with contextually demanding tasks. Overall, these results underscore the need for more inclusive training data and models better equipped to handle AAVE.
3.3 Intersection Over Union (IoU) Analysis
The IoU table shows the percentage of incorrectly answered questions in both our AAVE translation and SAE across five models and tasks. The results, shown in Tables 12 and 13 in the Appendix, indicate minimal overlap in incorrect responses, suggesting challenges in handling each dialect. The analysis reveals that challenges are dialect-specific, as there was minimal overlap in errors between SAE and AAVE. However, the high IoU in WSC for models like Gemini-1.5-Flash indicates that some tasks present had significant difficulties in both dialects. These results show the importance of developing more robust models that can handle the distinct features of AAVE, as current models show variability in managing dialect-specific errors.
4 Related Work
The development of benchmarks such as GLUE and SuperGLUE has significantly advanced the evaluation of language models on a variety of NLU tasks, including sentiment analysis, natural language inference, and reading comprehension (Wang et al., 2019, 2020). While these benchmarks have become standard tools for assessing model performance, their primary focus on SAE often leads to performance disparities when applied to non-standard dialects.
To address these disparities, the VALUE benchmark was introduced. VALUE incorporates deterministic linguistic transformations to evaluate the performance of the model in AAVE (Ziems et al., 2022, 2023). Although VALUE aims to provide a comprehensive evaluation of models processing dialectal variations, its deterministic nature can limit generalizability across various contexts, reducing broader application effectiveness.
Recent advances in LLM, such as GPT-4, have shown substantial improvements in NLU tasks, including sentiment analysis, machine translation, and question answering (Radford et al., 2019; Brown et al., 2020; OpenAI, 2024; Bubeck et al., 2023). Despite these advancements, research indicates that these models still exhibit biases against non-standard dialects like AAVE, particularly in tasks that involve natural language generation and toxicity detection (Zhou et al., 2021; Deas et al., 2023). This bias underscores the need for more inclusive benchmarks and evaluation frameworks.
The GEM benchmark and other studies have highlighted the importance of evaluating and mitigating biases in NLP models to promote fairness and inclusivity (Gehrmann et al., 2021; Moradi et al., 2024). These works emphasize the necessity of developing benchmarks that evaluate model performance across a range of dialects, ensuring that language technologies are equitable and reliable for all users.
By introducing AAVENUE, we aim to provide a comprehensive benchmark that evaluates LLM performance on both AAVE and SAE, promoting inclusivity in NLP tools and addressing the limitations of existing benchmarks. This approach aligns with recent research that emphasizes the need for more dialect-inclusive benchmarks to enhance the fairness and reliability of language technologies across diverse linguistic communities.
5 Conclusion
This paper introduced AAVENUE, a benchmark designed to evaluate LLMs on AAVE and SAE tasks. By leveraging GPT-4o-mini and few-shot prompting, we translated SAE tasks to AAVE. Our comprehensive evaluation, considering metrics like fluency, quality, and understandability, along with feedback from human validators, revealed that our translations were more superior then VALUES translations. Our evaluation results, revealed inherent biases in LLMs, highlighting a performance gap between SAE and AAVE. These findings build on the foundational work of others in this field, emphasizing the need for more inclusive training data and improved model architectures to address dialectal variations. We plan to extend this work by exploring additional dialects and further refining our translation methods. By doing so, we aim to set a new standard for equitable and accurate language processing across diverse communities.
6 Limitations
While AAVENUE provides a comprehensive benchmark for evaluating LLM performance across SAE and AAVE, it is not without its limitations. First, our dataset primarily focuses on a select number of tasks from the GLUE and SuperGLUE benchmarks, which may not fully capture the diversity of real-world applications where dialectal differences are prominent. Additionally, although our translations were validated by AAVE speakers, the inherent variability in AAVE usage across different regions and communities could introduce challenges in generalizing our findings. Furthermore, the reliance on GPT-4o-mini for translations, despite its advanced capabilities, may still reflect biases from its training data, potentially affecting the accuracy and fairness of the translations. Future work will need to address these limitations by expanding the dataset to include a broader range of tasks, incorporating a more diverse set of dialects, and exploring methods to reduce model bias.
Ethics Statement
We are mindful of the ethical implications of our research, which focuses on evaluating and addressing dialectal biases in LLMs through the development of the AAVENUE benchmark. While some data used in this study is publicly available, we also collected original data with careful consideration to ensure cultural and linguistic authenticity. The data collection process adhered to ethical guidelines, and all participants provided informed consent. Fluent AAVE speakers were recruited for the validation process to ensure the translations accurately reflect cultural and linguistic nuances. Throughout our research, we aimed to avoid potential harm and bias, with the goal of contributing to the development of more inclusive NLP systems. We have made efforts to report our findings transparently and objectively. We believe our research advances the field while adhering to rigorous ethical standards.
Appendix A Appendix
A.1 Few-Shot AAVE Translation from VALUE
Table 3: Sample Few-Shot Prompted AAVE Translations from VALUE
AAVE Translation
"I was bewildered, but I knew dat it was no gud asking his ass to explain."
"Cochran pontificated windily for da camera."
"I don’t want them to follow in my footsteps, as I ain’t go to no college, but I want them to go."
A.2 Detailed Comparison Scores
Table 4: LLM-Based Binary Preference Scores Across Tasks - GPT Models (AAVE/VALUE/About the Same)
Task GPT-4o-mini GPT-4-turbo GPT-4o
BoolQ (P) 88.79/10.33/0.88 94.51/4.62/0.88 94.95/4.84/0.22
BoolQ (Q) 82.33/17.14/0.53 78.13/20.08/1.58 80.34/18.72/0.84
MultiRC (P) 85.71/14.29/0.00 71.43/28.57/0.00 100.00/0.00/0.00
MultiRC (Q) 66.09/28.16/5.75 62.07/32.18/5.75 74.14/20.69/4.60
SST-2 87.25/9.92/2.83 80.88/12.04/7.08 87.39/9.92/2.69
COPA (P) 90.42/9.38/0.21 72.92/26.88/0.21 85.42/14.17/0.42
COPA (C1) 78.46/20.68/0.85 67.38/31.13/1.49 76.33/22.39/1.28
COPA (C2) 77.23/22.34/0.43 64.26/33.62/2.13 79.15/20.21/0.64
WSC (P) 83.80/16.20/0.00 73.99/26.01/0.00 86.76/13.08/0.16
Table 5: LLM-Based Binary Preference Scores Across Tasks - Gemini Models (AAVE/VALUE/About the Same)
Task Gemini-1.5-flash Gemini-1.5-pro
BoolQ (P) 93.85/6.15/0.00 88.13/10.99/0.88
BoolQ (Q) 75.18/24.40/0.42 44.16/53.42/2.42
MultiRC (P) 42.86/57.14/0.00 14.29/85.71/0.00
MultiRC (Q) 56.90/39.66/3.45 44.25/46.55/9.20
SST-2 84.70/12.89/2.41 72.24/20.40/7.37
COPA (P) 75.21/24.58/0.21 65.83/33.75/0.42
COPA (C1) 69.51/29.42/1.07 58.00/39.23/2.77
COPA (C2) 68.51/31.28/0.21 56.60/40.64/2.77
WSC (P) 84.11/15.26/0.62 57.17/42.21/0.62
A.3 Fluency Scores
Table 6: Fluency Scores Across Tasks
Task AAVE (GPT-4o-mini) VALUE
BoolQ (Passages) 70.7169 57.3199
BoolQ (Questions) 54.0808 52.8357
COPA (Choice 1) 58.8326 51.5955
COPA (Choice 2) 55.5880 54.7378
COPA (Premise) 54.9623 60.1386
MultiRC (Questions) 66.6923 59.2436
MultiRC (Paragraphs) 51.7319 57.6290
SST-2 (Sentences) 71.6592 64.7925
WSC (Paragraphs) 64.9487 63.2925
A.4 Coherence Scores
Table 7: Coherence Scores Across Tasks
Task AAVE (GPT-4o-mini) VALUE Translations
BoolQ (P) 74.3910 62.0983
BoolQ (Q) 56.2019 51.9285
MultiRC (P) 74.6919 65.1029
MultiRC (Q) 60.1021 56.1028
SST-2 83.1921 74.1029
COPA (P) 71.2058 63.5912
COPA (C1) 65.0852 54.5192
COPA (C2) 56.2019 51.2912
WSC (P) 64.0991 61.2017
A.5 Quality Scores
Table 8: Quality Scores Across Tasks
Task AAVE (GPT-4o-mini) VALUE Translations
BoolQ (P) 76.5699 58.2102
BoolQ (Q) 64.9691 56.7069
MultiRC (P) 64.9026 53.2962
MultiRC (Q) 62.7819 55.2493
SST-2 70.1046 66.9014
COPA (P) 67.6039 66.4782
COPA (C1) 73.5693 67.2891
COPA (C2) 75.2962 61.8783
WSC (P) 66.1021 66.3019
A.6 Understandability Scores
Table 9: Understandability Scores Across Tasks
Task AAVE (GPT-4o-mini) VALUE Translations
BoolQ (P) 76.5327 52.2594
BoolQ (Q) 64.4789 51.5689
MultiRC (P) 66.7032 57.1391
MultiRC (Q) 68.2280 53.2496
SST-2 82.5961 47.6228
COPA (P) 64.3259 57.6248
COPA (C1) 59.3256 48.1259
COPA (C2) 63.5832 57.3462
WSC (P) 76.7991 53.6917
A.7 BARTScores
Table 10: BARTScores Across Tasks
Task AAVE (GPT-4o-mini) VALUE Translations
BoolQ (P) -1.4402 -1.5352
BoolQ (Q) -1.6750 -2.8862
MultiRC (P) -1.8827 -1.7628
MultiRC (Q) -1.8490 -2.4334
SST-2 -2.5356 -2.5991
COPA (P) -1.8744 -3.0585
COPA (C1) -1.9161 -3.1870
COPA (C2) -2.0007 -3.1323
WSC (P) -1.9589 -2.5989
A.8 BLEU Scores
Table 11: BLEU Scores Across Tasks (Percentage Below Threshold)
Task < 0.7 < 0.5 < 0.3
BoolQ (Passage) 91.09% 57.66% 18.82%
BoolQ (Question) 79.38% 53.35% 35.64%
COPA (Premise) 87.20% 74.80% 59.40%
COPA (Choice 1) 85.40% 68.20% 58.60%
COPA (Choice 2) 80.60% 68.20% 56.40%
MultiRC (Paragraph) 100.00% 98.90% 95.00%
MultiRC (Question) 77.50% 55.20% 31.10%
SST-2 (Sentence) 96.60% 85.70% 64.10%
WSC (Paragraph) 88.15% 57.29% 23.71%
A.9 Intersection Over Union (IoU) Analysis
Table 12: Intersection Over Union Between Incorrect Answers for SAE and AAVE Across Tasks - GPT Models (%)
Task GPT-4o-mini GPT-4-turbo GPT-4o
SST-2 8.40 5.10 9.80
BoolQ 10.21 10.71 8.91
COPA 3.00 1.60 2.00
WSC 35.56 24.01 25.68
MultiRC 9.60 9.00 8.30
Table 13: Intersection Over Union Between Incorrect Answers for SAE and AAVE Across Tasks - Gemini Models (%)
Task Gemini-1.5-Flash Gemini-1.5-Pro
SST-2 10.40 6.20
BoolQ 8.51 8.41
COPA 5.80 1.80
WSC 49.54 44.53
MultiRC 9.90 7.90
A.10 Human Validators’ Scores
Table 14: Human Validators’ Average Scores for AAVE Translations Across COPA, BoolQ, and MultiRC Tasks (Out of 10)
Validator COPA BoolQ MultiRC
Validator 1 6.9 7.1 7.4
Validator 2 7.0 7.4 7.5
Validator 3 7.6 6.5 7.0
Validator 4 6.5 6.9 7.0
Validator 5 7.5 6.9 7.4
Validator 6 7.3 7.0 7.2
Validator 7 7.6 7.2 7.5
Validator 8 7.2 7.1 7.3
Validator 9 7.3 7.2 7.4
Validator 10 7.3 6.9 7.0
Average 7.22 7.02 7.27
Table 15: Human Validators’ Average Scores for AAVE Translations Across SST-2 and WSC Tasks (Out of 10)
Validator SST-2 WSC
Validator 1 6.6 7.3
Validator 2 7.4 8.3
Validator 3 7.2 6.9
Validator 4 6.6 7.0
Validator 5 7.0 6.9
Validator 6 7.3 7.2
Validator 7 7.5 7.1
Validator 8 7.1 7.4
Validator 9 7.3 7.3
Validator 10 6.9 7.1
Average 7.09 7.25

### v2

@@
 To address dialect-induced performance discrepancies, we introduce AAVENUE (AAVE Natural Language Understanding Evaluation), a benchmark for evaluating large language model (LLM) performance on NLU tasks in AAVE and Standard American English (SAE).
-AAVENUE builds upon and extends existing benchmarks like VALUE, replacing deterministic syntactic and morphological transformations with a more flexible methodology leveraging LLM-based translation with few-shot prompting, improving performance across our evaluation metrics when translating key tasks from the GLUE and SuperGLUE benchmarks.
+AAVENUE builds upon and extends existing benchmarks like VALUE, replacing deterministic syntactic and morphological transformations with a more flexible methodology leveraging LLM-based translation with few-shot prompting, improving performance across several evaluation metrics when translating key tasks from the GLUE and SuperGLUE benchmarks.
 We compare AAVENUE and VALUE translations using five popular LLMs and a comprehensive set of metrics including fluency, BARTScore, quality, coherence, and understandability.
-Additionally, we recruit fluent AAVE speakers to validate our translations for authenticity.
+Additionally, the fluency of AAVENUE is validated by annotations from AAVE speakers.
 Our evaluations reveal that LLMs consistently perform better on SAE tasks than AAVE-translated versions, underscoring inherent biases and highlighting the need for more inclusive NLP models.
+This work has been accepted at the NLP4I workshop at EMNLP 2024.
 We have open-sourced our source code on GitHub and created a website to showcase our work at https://aavenue.live.
 AAVENUE: Detecting LLM Biases on NLU Tasks in AAVE via a Novel Benchmark
-Abhay Guptathanks: Lead Author           Philip Meng           Ece Yurtseven           Sean O’Brien thanks: Senior Author          Kevin Zhu22footnotemark: 2 Algoverse AI Research abhaygupta1266@gmail.com, kevin@algoverse.us
+Abhay Guptathanks: Lead Author           Ece Yurtseven           Philip Meng           Sean O’Brienthanks: Senior Authors           Kevin Zhu22footnotemark: 2 Algoverse AI Research abhaygupta1266@gmail.com, kevin@algoverse.us
 1 Introduction
@@
 2 Dataset
-To evaluate the performance of models on SAE and AAVE, we selected five key tasks from the GLUE and SuperGLUE benchmarks, knwon for their diversity and relevance to natural language understanding tasks (Wang et al., 2019, 2020).
+To evaluate the performance of models on SAE and AAVE, we selected five key tasks from the GLUE and SuperGLUE benchmarks, known for their diversity and relevance to natural language understanding tasks (Wang et al., 2019, 2020).
 These tasks include BoolQ, MultiRC, SST-2, COPA, and WSC.
@@
 To see the few-shot prompt translations we used, please refer to Table 5 in the Appendix.
+Task Quality Fluency Coherence Understandability BARTScore
+AAVE VALUE AAVE VALUE AAVE VALUE AAVE VALUE AAVE VALUE
+BoolQ (P) 76.57 58.21 70.72 57.32 74.39 62.10 76.53 52.26 -1.44 -1.54
+BoolQ (Q) 64.97 56.71 54.08 52.84 56.20 51.93 64.48 51.57 -1.68 -2.89
+MultiRC (P) 64.90 53.30 51.73 57.63 74.69 65.10 66.70 57.14 -1.88 -1.76
+MultiRC (Q) 62.78 55.25 66.69 59.24 60.10 56.10 68.23 53.25 -1.85 -2.43
+SST-2 70.10 66.90 71.66 64.79 83.19 74.10 82.60 47.62 -2.54 -2.60
+COPA (P) 67.60 66.48 54.96 60.14 71.21 63.59 64.33 57.62 -1.87 -3.06
+COPA (C1) 73.57 67.29 58.83 51.60 65.09 54.52 59.33 48.13 -1.92 -3.19
+COPA (C2) 75.30 61.88 55.59 54.74 56.20 51.29 63.58 57.35 -2.00 -3.13
+WSC (P) 66.10 66.30 64.95 63.29 64.10 61.20 76.80 53.69 -1.96 -2.60
+Table 1: Comparison of Translation Metrics Across Tasks
+Task GPT-4o-mini GPT-4-turbo GPT-4o Gemini-1.5-Flash Gemini-1.5-Pro
+BoolQ (P) 88.79/10.33/0.88 94.51/4.62/0.88 94.95/4.84/0.22 93.85/6.15/0.00 88.13/10.99/0.88
+BoolQ (Q) 82.33/17.14/0.53 78.13/20.08/1.58 80.34/18.72/0.84 75.18/24.40/0.42 44.16/53.42/2.42
+MultiRC (P) 85.71/14.29/0.00 71.43/28.57/0.00 100.00/0.00/0.00 42.86/57.14/0.00 14.29/85.71/0.00
+MultiRC (Q) 66.09/28.16/5.75 62.07/32.18/5.75 74.14/20.69/4.60 56.90/39.66/3.45 44.25/46.55/9.20
+SST-2 87.25/9.92/2.83 80.88/12.04/7.08 87.39/9.92/2.69 84.70/12.89/2.41 72.24/20.40/7.37
+COPA (P) 90.42/9.38/0.21 72.92/26.88/0.21 85.42/14.17/0.42 75.21/24.58/0.21 65.83/33.75/0.42
+COPA (C1) 78.46/20.68/0.85 67.38/31.13/1.49 76.33/22.39/1.28 69.51/29.42/1.07 58.00/39.23/2.77
+COPA (C2) 77.23/22.34/0.43 64.26/33.62/2.13 79.15/20.21/0.64 68.51/31.28/0.21 56.60/40.64/2.77
+WSC (P) 83.80/16.20/0.00 73.99/26.01/0.00 86.76/13.08/0.16 84.11/15.26/0.62 57.17/42.21/0.62
+Table 2: LLM-Based Binary Preference Scores Across Tasks for GPT and Gemini Models (AAVE/VALUE/About the Same)
 2.3 Validation Steps
@@
 This shows that our translations are better in terms of overall quality, including accuracy, style, and appropriateness.
-(see Table 8).
 (2) Fluency: Our translations achieved a fluency score of 70.72 in BoolQ passages, compared to VALUE’s 57.32.
 This suggests that our translations are better written and more grammatically sound, resulting in improved readability.
-(see Table 6).
-(3) Coherence: Our translations exhibited better coherence, with scores like 74.39 for BoolQ passages versus VALUE’s 62.10, showing that our translations make more logical sense and maintain consistency throughout the text (see Table 7).
+(3) Coherence: Our translations exhibited better coherence, with scores like 74.39 for BoolQ passages versus VALUE’s 62.10, showing that our translations make more logical sense and maintain consistency throughout the text.
 (4) Understandability: In terms of understandability, our translations were clearer and more accessible, scoring 76.53 in BoolQ passages compared to 52.26 for VALUE.
-This indicates that our translations are easier for readers to understand (see Table 9).
-(5) BARTScores: Finally, our translations achieved better BARTScores, with a score of -1.44 in BoolQ passages compared to VALUE’s -1.54, indicating that our translations are closer to human-produced texts and better maintain fidelity to the original content (see Table 10).
+This indicates that our translations are easier for readers to understand.
+(5) BARTScores: Finally, our translations achieved better BARTScores, with a score of -1.44 in BoolQ passages compared to VALUE’s -1.54, indicating that our translations are closer to human-produced texts and better maintain fidelity to the original content.
 Collectively, these metrics confirm that our translations consistently outperform those from the VALUE benchmark, providing superior quality, fluency, coherence, understandability, and fidelity.
+Task GPT-4o-mini GPT-4-turbo GPT-4o Gemini-1.5-Flash Gemini-1.5-Pro
+SST-2 90.40/88.40 (-2.0) 94.00/92.80 (-1.2) 88.80/87.30 (-1.5) 87.70/87.10 (-0.6) 92.00/91.40 (-0.6)
+BoolQ 88.29/85.29 (-3.0) 88.09/86.49 (-1.6) 89.19/86.89 (-2.3) 89.69/87.29 (-2.4) 89.49/85.89 (-3.6)
+COPA 95.40/93.20 (-2.2) 97.60/96.80 (-0.8) 97.20/96.40 (-0.8) 91.40/92.00 (+0.6) 97.40/95.80 (-1.6)
+WSC 60.03/57.90 (-2.1) 69.60/68.69 (-0.9) 70.36/67.02 (-3.3) 48.78/48.48 (-0.3) 51.37/51.22 (-0.2)
+MultiRC 84.50/72.00 (-12.5) 86.20/73.70 (-12.5) 87.50/71.30 (-16.2) 84.10/70.70 (-13.4) 85.90/71.90 (-14.0)
+Table 3: Accuracy Scores for GPT and Gemini Models (%) All scores are presented in the format SAE/AAVE.
+Task GPT-4o-mini GPT-4-turbo GPT-4o Gemini-1.5-Flash Gemini-1.5-Pro
+SST-2 8.40 5.10 9.80 10.40 6.20
+BoolQ 10.21 10.71 8.91 8.51 8.41
+COPA 3.00 1.60 2.00 5.80 1.80
+WSC 35.56 24.01 25.68 49.54 44.53
+MultiRC 9.60 9.00 8.30 9.90 7.90
+Table 4: Intersection Over Union Between Incorrect Answers for SAE and AAVE Across Tasks (%)
 2.3.2 Comparison Scores Analysis
 The comparison scores provide a direct evaluation of our AAVE translations against those from the VALUE benchmark across various tasks and models.
-As shown in Tables 4 and 5, our AAVE translations were consistently preferred over the VALUE translations by the LLMs we evaluated.
-For instance, in the BoolQ passage task using GPT-4-turbo, our translations were preferred 94.51% of the time compared to VALUE’s 4.62% (see Table 4).
-Similarly, in the COPA premise task, GPT-4o-mini showed a preference for our translations 90.42% of the time, with VALUE translations receiving only 9.38% (see Table 4).
-When evaluated with the Gemini-1.5-flash model on the BoolQ passage task, our translations were preferred 93.85% of the time compared to VALUE’s 6.15% (see Table 5).
-These consistent preferences across multiple models and tasks show the effectiveness of our translations in capturing the nuances of AAVE while maintaining the meaning of the original text.
+As shown in Table 2, our AAVE translations were consistently preferred over the VALUE translations by the LLMs we evaluated.
+For instance, in the BoolQ passage task using GPT-4-turbo, our translations were preferred 94.51% of the time compared to VALUE’s 4.62%.
+Similarly, in the COPA premise task, GPT-4o-mini showed a preference for our translations 90.42% of the time, with VALUE translations receiving only 9.38%.
+When evaluated with the Gemini-1.5-Flash model on the BoolQ passage task, our translations were preferred 93.85% of the time compared to VALUE’s 6.15%.
+These consistent preferences across multiple models and tasks demonstrate the effectiveness of our translations in capturing the nuances of AAVE while maintaining the meaning of the original text.
 The results indicate that our approach to translating SAE to AAVE not only meets but often exceeds the standards set by the VALUE benchmark.
@@
 We recruited 10 fluent AAVE speakers from the Bronx and Queens area to assess the quality of our AAVE translations across five tasks on a scale of 1 to 10, specifically rating how accurately the translations reflect AAVE.
-The results, detailed in Tables 14 and 15, show that the average scores across the tasks ranged from 7.02 to 7.27.
+The results, detailed in Table 6 in Section B, show that the average scores across the tasks ranged from 7.02 to 7.27.
 The highest average score was observed for MultiRC (7.27), while BoolQ received the lowest average score (7.02).
-These consistent ratings across multiple fluent speakers, reinforce the reliability of our AAVE translations.
-The slightly higher scores for subsets like WSC and MultiRC suggest that our translations were particularly effective in maintaining clarity and coherence in more complex linguistic structures.
+These consistent ratings across multiple fluent speakers reinforce the reliability of our AAVE translations.
+The slightly higher scores for tasks like WSC and MultiRC suggest that our translations were particularly effective in maintaining clarity and coherence in more complex linguistic structures.
 Overall, the validators’ feedback confirms the quality and authenticity of our translations, aligning well with our evaluation metrics and further validating our approach.
 3 Results
-3.1 Accuracy Scores Across Tasks
-We evaluated the accuracy of the translations in five tasks using five LLMs. The accuracy scores show the performance of each model in SAE and AAVE translations, as shown in Tables 1 and 2.
-Table 1: Accuracy Scores for GPT Models (%)
-All scores are presented in the format SAE/AAVE.
-Task GPT-4o-mini GPT-4-turbo GPT-4o
-SST-2 90.40/88.40 (-2.0) 94.00/92.80 (-1.2) 88.80/87.30 (-1.5)
-BoolQ 88.29/85.29 (-3.0) 88.09/86.49 (-1.6) 89.19/86.89 (-2.3)
-COPA 95.40/93.20 (-2.2) 97.60/96.80 (-0.8) 97.20/96.40 (-0.8)
-WSC 60.03/57.90 (-2.1) 69.60/68.69 (-0.9) 70.36/67.02 (-3.3)
-MultiRC 84.50/72.00 (-12.5) 86.20/73.70 (-12.5) 87.50/71.30 (-16.2)
-Table 2: Accuracy Scores for Gemini Models (%)
-All scores are presented in the format SAE/AAVE.
-Task Gemini-1.5-Flash Gemini-1.5-Pro
-SST-2 87.70/87.10 (-0.6) 92.00/91.40 (-0.6)
-BoolQ 89.69/87.29 (-2.4) 89.49/85.89 (-3.6)
-COPA 91.40/92.00 (+0.6) 97.40/95.80 (-1.6)
-WSC 48.78/48.48 (-0.3) 51.37/51.22 (-0.2)
-MultiRC 84.10/70.70 (-13.4) 85.90/71.90 (-14.0)
+We evaluated the accuracy of the translations in five tasks using five LLMs.
+The accuracy scores show the performance of each model in SAE and AAVE translations, highlighting consistent performance drops when handling AAVE translations across all models, as shown in Table 3.
 3.1 Accuracy Score Analysis
-The accuracy scores from Tables 1 and 2 highlight consistent performance drops when handling AAVE translations across all models.
+The accuracy scores from Table 3 highlight consistent performance drops when handling AAVE translations across all models.
 MultiRC and WSC exhibited the largest declines, indicating challenges in complex reading comprehension and pronoun resolution tasks.
@@
 The IoU table shows the percentage of incorrectly answered questions in both our AAVE translation and SAE across five models and tasks.
-The results, shown in Tables 12 and 13 in the Appendix, indicate minimal overlap in incorrect responses, suggesting challenges in handling each dialect.
+The results, shown in Table 4, indicate minimal overlap in incorrect responses, suggesting challenges in handling each dialect.
 The analysis reveals that challenges are dialect-specific, as there was minimal overlap in errors between SAE and AAVE.
@@
 The data collection process adhered to ethical guidelines, and all participants provided informed consent.
-Fluent AAVE speakers were recruited for the validation process to ensure the translations accurately reflect cultural and linguistic nuances.
+Our human validators, who are fluent AAVE speakers from the Bronx, NY, were recruited to ensure the translations accurately reflect cultural and linguistic nuances.
+Annotators were compensated for their time and effort, and we encouraged them to take breaks if they felt overwhelmed during the annotation process.
 Throughout our research, we aimed to avoid potential harm and bias, with the goal of contributing to the development of more inclusive NLP systems.
@@
 We believe our research advances the field while adhering to rigorous ethical standards.
-Appendix A Appendix
-A.1 Few-Shot AAVE Translation from VALUE
-Table 3: Sample Few-Shot Prompted AAVE Translations from VALUE
-AAVE Translation
+Appendix A Few-Shot AAVE Translation Examples
+AAVE Translation Examples
 "I was bewildered, but I knew dat it was no gud asking his ass to explain."
@@
 "I don’t want them to follow in my footsteps, as I ain’t go to no college, but I want them to go."
-A.2 Detailed Comparison Scores
-Table 4: LLM-Based Binary Preference Scores Across Tasks - GPT Models (AAVE/VALUE/About the Same)
-Task GPT-4o-mini GPT-4-turbo GPT-4o
-BoolQ (P) 88.79/10.33/0.88 94.51/4.62/0.88 94.95/4.84/0.22
-BoolQ (Q) 82.33/17.14/0.53 78.13/20.08/1.58 80.34/18.72/0.84
-MultiRC (P) 85.71/14.29/0.00 71.43/28.57/0.00 100.00/0.00/0.00
-MultiRC (Q) 66.09/28.16/5.75 62.07/32.18/5.75 74.14/20.69/4.60
-SST-2 87.25/9.92/2.83 80.88/12.04/7.08 87.39/9.92/2.69
-COPA (P) 90.42/9.38/0.21 72.92/26.88/0.21 85.42/14.17/0.42
-COPA (C1) 78.46/20.68/0.85 67.38/31.13/1.49 76.33/22.39/1.28
-COPA (C2) 77.23/22.34/0.43 64.26/33.62/2.13 79.15/20.21/0.64
-WSC (P) 83.80/16.20/0.00 73.99/26.01/0.00 86.76/13.08/0.16
-Table 5: LLM-Based Binary Preference Scores Across Tasks - Gemini Models (AAVE/VALUE/About the Same)
-Task Gemini-1.5-flash Gemini-1.5-pro
-BoolQ (P) 93.85/6.15/0.00 88.13/10.99/0.88
-BoolQ (Q) 75.18/24.40/0.42 44.16/53.42/2.42
-MultiRC (P) 42.86/57.14/0.00 14.29/85.71/0.00
-MultiRC (Q) 56.90/39.66/3.45 44.25/46.55/9.20
-SST-2 84.70/12.89/2.41 72.24/20.40/7.37
-COPA (P) 75.21/24.58/0.21 65.83/33.75/0.42
-COPA (C1) 69.51/29.42/1.07 58.00/39.23/2.77
-COPA (C2) 68.51/31.28/0.21 56.60/40.64/2.77
-WSC (P) 84.11/15.26/0.62 57.17/42.21/0.62
-A.3 Fluency Scores
-Table 6: Fluency Scores Across Tasks
-Task AAVE (GPT-4o-mini) VALUE
-BoolQ (Passages) 70.7169 57.3199
-BoolQ (Questions) 54.0808 52.8357
-COPA (Choice 1) 58.8326 51.5955
-COPA (Choice 2) 55.5880 54.7378
-COPA (Premise) 54.9623 60.1386
-MultiRC (Questions) 66.6923 59.2436
-MultiRC (Paragraphs) 51.7319 57.6290
-SST-2 (Sentences) 71.6592 64.7925
-WSC (Paragraphs) 64.9487 63.2925
-A.4 Coherence Scores
-Table 7: Coherence Scores Across Tasks
-Task AAVE (GPT-4o-mini) VALUE Translations
-BoolQ (P) 74.3910 62.0983
-BoolQ (Q) 56.2019 51.9285
-MultiRC (P) 74.6919 65.1029
-MultiRC (Q) 60.1021 56.1028
-SST-2 83.1921 74.1029
-COPA (P) 71.2058 63.5912
-COPA (C1) 65.0852 54.5192
-COPA (C2) 56.2019 51.2912
-WSC (P) 64.0991 61.2017
-A.5 Quality Scores
-Table 8: Quality Scores Across Tasks
-Task AAVE (GPT-4o-mini) VALUE Translations
-BoolQ (P) 76.5699 58.2102
-BoolQ (Q) 64.9691 56.7069
-MultiRC (P) 64.9026 53.2962
-MultiRC (Q) 62.7819 55.2493
-SST-2 70.1046 66.9014
-COPA (P) 67.6039 66.4782
-COPA (C1) 73.5693 67.2891
-COPA (C2) 75.2962 61.8783
-WSC (P) 66.1021 66.3019
-A.6 Understandability Scores
-Table 9: Understandability Scores Across Tasks
-Task AAVE (GPT-4o-mini) VALUE Translations
-BoolQ (P) 76.5327 52.2594
-BoolQ (Q) 64.4789 51.5689
-MultiRC (P) 66.7032 57.1391
-MultiRC (Q) 68.2280 53.2496
-SST-2 82.5961 47.6228
-COPA (P) 64.3259 57.6248
-COPA (C1) 59.3256 48.1259
-COPA (C2) 63.5832 57.3462
-WSC (P) 76.7991 53.6917
-A.7 BARTScores
-Table 10: BARTScores Across Tasks
-Task AAVE (GPT-4o-mini) VALUE Translations
-BoolQ (P) -1.4402 -1.5352
-BoolQ (Q) -1.6750 -2.8862
-MultiRC (P) -1.8827 -1.7628
-MultiRC (Q) -1.8490 -2.4334
-SST-2 -2.5356 -2.5991
-COPA (P) -1.8744 -3.0585
-COPA (C1) -1.9161 -3.1870
-COPA (C2) -2.0007 -3.1323
-WSC (P) -1.9589 -2.5989
-A.8 BLEU Scores
-Table 11: BLEU Scores Across Tasks (Percentage Below Threshold)
-Task < 0.7 < 0.5 < 0.3
+Table 5: Examples of the Few-Shot Prompted AAVE Translations from VALUE used in our experiments.
+Appendix B Human Validators’ Scores
+Validator COPA BoolQ MultiRC SST-2 WSC
+Validator 1 6.9 7.1 7.4 6.6 7.3
+Validator 2 7.0 7.4 7.5 7.4 8.3
+Validator 3 7.6 6.5 7.0 7.2 6.9
+Validator 4 6.5 6.9 7.0 6.6 7.0
+Validator 5 7.5 6.9 7.4 7.0 6.9
+Validator 6 7.3 7.0 7.2 7.3 7.2
+Validator 7 7.6 7.2 7.5 7.5 7.1
+Validator 8 7.2 7.1 7.3 7.1 7.4
+Validator 9 7.3 7.2 7.4 7.3 7.3
+Validator 10 7.3 6.9 7.0 6.9 7.1
+Average 7.22 7.02 7.27 7.09 7.25
+Table 6: Human Validators’ Scores for AAVE Translations Across All Tasks (Out of 10)
+Appendix C BLEU Scores
+Task BLEU <<< 0.7 (%) BLEU <<< 0.5 (%) BLEU <<< 0.3 (%)
 BoolQ (Passage) 91.09% 57.66% 18.82%
@@
 WSC (Paragraph) 88.15% 57.29% 23.71%
-A.9 Intersection Over Union (IoU) Analysis
-Table 12: Intersection Over Union Between Incorrect Answers for SAE and AAVE Across Tasks - GPT Models (%)
-Task GPT-4o-mini GPT-4-turbo GPT-4o
-SST-2 8.40 5.10 9.80
-BoolQ 10.21 10.71 8.91
-COPA 3.00 1.60 2.00
-WSC 35.56 24.01 25.68
-MultiRC 9.60 9.00 8.30
-Table 13: Intersection Over Union Between Incorrect Answers for SAE and AAVE Across Tasks - Gemini Models (%)
-Task Gemini-1.5-Flash Gemini-1.5-Pro
-SST-2 10.40 6.20
-BoolQ 8.51 8.41
-COPA 5.80 1.80
-WSC 49.54 44.53
-MultiRC 9.90 7.90
-A.10 Human Validators’ Scores
-Table 14: Human Validators’ Average Scores for AAVE Translations Across COPA, BoolQ, and MultiRC Tasks (Out of 10)
-Validator COPA BoolQ MultiRC
-Validator 1 6.9 7.1 7.4
-Validator 2 7.0 7.4 7.5
-Validator 3 7.6 6.5 7.0
-Validator 4 6.5 6.9 7.0
-Validator 5 7.5 6.9 7.4
-Validator 6 7.3 7.0 7.2
-Validator 7 7.6 7.2 7.5
-Validator 8 7.2 7.1 7.3
-Validator 9 7.3 7.2 7.4
-Validator 10 7.3 6.9 7.0
-Average 7.22 7.02 7.27
-Table 15: Human Validators’ Average Scores for AAVE Translations Across SST-2 and WSC Tasks (Out of 10)
-Validator SST-2 WSC
-Validator 1 6.6 7.3
-Validator 2 7.4 8.3
-Validator 3 7.2 6.9
-Validator 4 6.6 7.0
-Validator 5 7.0 6.9
-Validator 6 7.3 7.2
-Validator 7 7.5 7.1
-Validator 8 7.1 7.4
-Validator 9 7.3 7.3
-Validator 10 6.9 7.1
-Average 7.09 7.25
+Table 7: BLEU Scores Across Tasks (Percentage of Translations Below BLEU Thresholds). These scores measure the lexical similarity of our translations to the original texts.